A platinium resistance thermometer has resistance of 52.5 ohms and 9.75 ohms at 0 degrees Celsius and 100 degrees Celsius respectively. When the resistance is 8.25 ohms, find the temperature.
1
Expert's answer
2012-04-10T07:19:37-0400
Let's find out the relation between resistance and temperature. We assume that resistance changes linearly with respect to temperature, i.e.
R = αT.
Let's find coefficient α.
Difference in temperature: 100-0 = 100 C.
Difference in resistance: 52.5-9.75 = 42.75 Ohms.
Therefore,
& alpha; = 42.75/100 = 0.4275 Ohm/K.
So, the temperature corresponding to the resistance of 8.25 ohms is
T = R/α = 8.25/0.4275 ≈ 19.2982 K.